A first number plus twice a second number is 8. Twice the first number, plus the second number totals 25. Find the numbers.
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

x+2y=8 .............1
2x+y=25 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -2
1 x + 2 y = 8
-4 x + -2 y = -50
Add the two equations
-3 x = -42
/ -3
x=14
plug value of x in (1)
1x+2y=8
14+2y=8
2y=8-14
2y=-6
y=-3

Answer by ikleyn(53749)

.
A first number plus twice a second number is 8. Twice the first number, plus the second number totals 25.
Find the numbers.
~~~~~~~~~~~~~~~~~~~~~~~~


        There is very easy and effective way to solve it - practically mentally,  in the head.


The equations are

     x + 2y =  8    (1)

    2x +  y = 25    (2)


Add the equations

    3x + 3y = 33,

hence

    x + y = 11.     (3).


Compare (1) and (3).  It should be clear from the comparison that  y = -3.


Next, compare equations (2) and (3).  It should be clear for the comparison that  x = 14.


Now check these values and makes sure that they satisfy equations.

That's all.